The base of a triangle sail 8m shorter than its height the area is 24m^2. find the base and length

Answer by checkley79(3341)

A=BH/2
B=H-8
24=(H-8)H/2
24=(H^2-8H)/2 CROSS MULTIPLY
H^2-8H=24*2
H^2-8H=48
H^2-8H-48=0
(H-12)(H+4)=0
H-12=O
H=12 ANS. FOR THE HEIGHT.
B=12-8=4 ANS. FOR THE BASE.
PROOF:
24=12*4/2
24=48/2
24=24